### v1

# Uni-ISP: Unifying the Learning of ISPs from Multiple Cameras

Abstract
Modern end-to-end image signal processors (ISPs) can learn complex mappings from RAW/XYZ data to sRGB (or inverse), opening new possibilities in image processing. However, as the diversity of camera models continues to expand, developing and maintaining individual ISPs is not sustainable in the long term, which inherently lacks versatility, hindering the adaptability to multiple camera models. In this paper, we propose a novel pipeline, Uni-ISP, which unifies the learning of ISPs from multiple cameras, offering an accurate and versatile processor to multiple camera models. The core of Uni-ISP is leveraging device-aware embeddings through learning inverse/forward ISPs and its special training scheme. By doing so, Uni-ISP not only improves the performance of inverse/forward ISPs but also unlocks a variety of new applications inaccessible to existing learned ISPs. Moreover, since there is no dataset synchronously captured by multiple cameras for training, we construct a real-world 4K dataset, FiveCam, comprising more than 2,400 pairs of sRGB-RAW images synchronously captured by five smartphones. We conducted extensive experiments demonstrating Uni-ISP’s accuracy in inverse/forward ISPs (with improvements of +1.5dB/2.4dB PSNR), its versatility in enabling new applications, and its adaptability to new camera models.
Figure 1: We propose Uni-ISP, a model that unifies the learning of inverse and forward ISP behaviors of multiple cameras simultaneously. By leveraging the shared characteristics across various camera ISPs, our method can achieve higher performance in inverse and forward ISP (A) compared to previously learned ISP methods tailored for only one camera separately. Meanwhile, the device-aware property of the Uni-ISP enables new cross-camera ISP applications for a learned ISP model, including photographic appearance transfer (B and C), inter/extrapolation (D), and zero-shot image forensics (E and F).
1 Introduction
Image Signal Processor (ISP) transforms raw image data captured by camera sensors into viewable formats such as sRGB, playing a pivotal role in determining the visual quality of photographs [16]. Through the meticulous design of their ISPs, various camera brands have cultivated distinctive photographic appearances that resonate with diverse user preferences [9, 33]. For example, Apple’s smartphone cameras are celebrated for their sharp and distinctive Apple feel, while Leica cameras are esteemed for their glow and deep color tones, contributing to the iconic Leica look.
Recently, neural networks have been used to approximate the entire ISP or a specific module, aka learned ISP, bringing two main benefits. 1) Performance Enhancement. The powerful representation capabilities of neural networks enable learned ISPs to perform challenging tasks, such as hallucinating detailed content in highlights and shadows [42]. 2) New Functionalities. Learned ISPs introduce new functionalities, such as inverse ISP [38, 1], which converts sRGB images back to RAW/XYZ space, offering greater flexibility and potential for raw-domain enhancement and further manipulation, such as deblurring [22], denoising [32, 5, 3], HDR photography [23, 42], etc. These innovations expand the applications and potential of learned ISPs. However, current methods still design and train ISPs for individual camera models, which might limit the synergistic benefits across different ISPs (see Section 4.1). Moreover, versatile models have shown advantages in various fields in low-level vision [31, 21], high-level vision [19], and multi-modality [41, 39, 35]. As the number of camera models grows, individual learned ISPs may also lack the versatility and adaptability for widespread use, potentially making it unsustainable in the long term.
In this paper, we aim to unify the learning of ISPs from different cameras, which offers two direct advantages. First, it enhances visual quality by leveraging the synergies among ISPs from multiple cameras. Unified learning enables ISPs to understand the underlying commonalities and differences in the diverse data, leading to overall improved visual performance. Second, unified learning provides novel applications, surpassing the capabilities of existing learned ISPs limited to inverse and forward ISPs. The unified learning enables new uses such as photographic appearance transfer, interpolation, and extrapolation across diverse camera models. Additionally, it supports zero-shot image forensics based on the self-consistency of ISP behaviors, including image-level source camera identification and pixel-level image splice detection.
However, learning a device-aware ISP model for multiple cameras is far from trivial, presenting several challenges. First, we observe that simply mixing training data from multiple cameras in current models [1, 40, 38, 18] can not bring satisfactory performances. Therefore, we propose Uni-ISP, a novel unified ISP model for multiple cameras that contains several optimizable device-aware embeddings to learn the ISPs of different cameras. These device-aware embeddings enable the model to capture the specific characteristics tailored to individual devices, while the shared backbones capture underlying commonalities.
Second, current ISP datasets do not contain synchronized sRGB-Raw image pairs captured by multiple cameras. Although this data is not necessary to learn the individual ISP, it is critical for learning a unified ISP that supports synergistic benefits across different camera models and enables the development of new applications. To address this, we develop a synchronized camera array with five smartphones and construct a novel dataset, FiveCam, which consists of 2,464 synchronized high-quality sRGB-Raw paired images with 4K spatial resolution. The captured dataset has a wide range of scenarios, from landscape to close-up, and contains different lighting conditions, including both indoor and outdoor settings during day and night.
Third, given the inevitable misalignment in synchronized photo pairs taken by different cameras, a robust alignment and training scheme is required. To tackle this challenge, we first roughly align the images using optical-flow-based methods, which will introduce frequency bias in wrapped images. Then, we design a frequency bias correction (FBC) loss to mitigate texture blur. Additionally, we introduce the self-/cross-camera training schemes to facilitate applications across the same/different camera models.
With all these three designs, our Uni-ISP can be applied to a wide range of image tasks (Figure 1), such as photographic appearance transfer, interpolation, and extrapolation across diverse camera models. Users can apply the visual characteristics of one camera model to another, achieving unique aesthetic effects. It also facilitates zero-shot image forensics tasks by utilizing the self-consistence of these ISP behaviors, including source camera identification and image splice detection. Extensive experiments show that Uni-ISP outperforms state-of-the-art methods with approximately 1.5dB PSNR in inverse ISP and approximately 2.4dB PSNR in forward ISP.
Figure 2: The model design of Uni-ISP. Uni-ISP contains two modules, the inverse ISP module $g$ and the forward ISP module $h$. Both two modules share the same structure. For visual simplicity, we draw the inverse ISP module $g$ as a thumbnail, whose inner structure is the same as the forward ISP module $h$. The device-aware embeddings are optimizable parameters and will be selected to interact with the bottleneck features via the DEIM during the training or inference.
2 Related Work
Learning Inverse and Forward ISP
Training neural networks to learn the inverse and forward processes of certain ISPs has been explored in recent research works [1, 40, 38]. CycleISP [40] maintains the self-contained property of inverse and forward ISP, where the cycle constraints are applied in addition to the independent supervision from the RGB or the raw ground truth. The InvISP [38] achieves the inverse and forward ISP in one self-contained invertible network. However, constraints that ensure the model’s invertibility also limit the network’s expression ability. ParamISP [18] utilizes the EXIF data in the model design. All existing methods ignore the variance of camera devices, which overlooks the potential value of the commonalities across different camera devices in the learned ISP task. In contrast, our Uni-ISP leverages these underlying commonalities by unifying learning of multiple cameras.
Photorealistic Image Style Transfer
Photorealistic image style transfer manipulates the aesthetic of an image, such as its color and tone, without distorting its original structure and content. Non-learning-based photorealistic image style transfer is primarily handled through traditional image processing techniques [29, 30, 27], including color grading [30], gamut optimization [27], etc.  Gharbi et al. [13] and  Xia et al. [37] propose learning-based bilateral grids for photorealistic image style transfer. The recent large generative models have also opened up new possibilities for photorealistic image style transfer [12, 20, 24, 4], but at a cost of huge computational resources and the risk of distorting the original content. For efficient and high-fidelity photorealistic image style transfer, NeuralPreset [17] is proposed for learning photorealistic image style transfer with a small memory footprint, which can be easily deployed on mobile devices. Unlike the aforementioned methods, our Uni-ISP learns the transfer of style-like visual feeling produced by certain camera ISPs, the photographic appearance, which is device-dependent and physically faithful.
Image Forensics
Image forensics focuses on validating and analyzing digital images to determine their origin by source camera identification or authenticity by splice detection. Source camera identification is highly related to the properties of the camera, which can be done by comparing the sensor pattern noise (SPN) and its main component photo response non-uniformity (PRNU) noise of a given image taken with an unknown camera and the reference image taken with known cameras.  Lukas et al. [25] extract the photo response non-uniformity (PRNU) using the discrete wavelet transform.  Chen et al. [7] propose a residual network for source camera identification.  Hui et al. [14] propose a multi-scale feature fusion network and a corresponding two-stage training scheme for this task based on the guidance of PRNU. Some methods are not based on the PRNU but on the lens distortions and auto white balance (AWB) algorithms [10, 2]. Unlike the existing methods that require specific training, our method achieves zero-shot source camera identification by implicitly identifying the whole ISP behavior. Our model can also perform image splice detection by slightly modifying the inference process for source camera identification.
3 Method
3.1 Overview
First, we discuss the XYZ image format of the inverse and forward ISP tasks that our model will undertake. The XYZ images are device-independent measurements of radiances and learning XYZ images enjoys the same benefits as learning raw images. Therefore, in line with [1], we opt for XYZ images processed from the raw images taken by the camera as the raw modality. Specifically, the XYZ image is obtained from the real raw image by applying the early fixed stage of the ISP using the as-shot white balance, a fixed linear demosaicing algorithm, and the camera-to-XYZ matrix of the current device, without applying gamma tone mapping. In this setting, the XYZ images are linearly correlated to the raw images and they can be converted to each other without loss.
Figure 2 shows the overview of Uni-ISP, which contains the inverse ISP module $g$ and forward ISP module $h$. Our model aims to be aware of various camera devices in the learned ISP tasks. Suppose the images we discussed here are all in size of $H\times W$ and channels of $C$. Given an sRGB image $I_{a}\in\mathbb{R}^{H\times W\times C}$ produced by the camera $a$, $\mathcal{E}_{a}\in\mathbb{R}^{D}$ represents the device-aware embedding for camera $a$ with size of $D$, the inverse ISP module $g$ learns to output the corresponding XYZ image $\hat{L}_{a}$ of camera $a$ given the input $I_{a}$:
$\hat{L}_{a}=g(I_{a},\mathcal{E}_{a}),$ (1)
and forward ISP module $h$ learns to predict the $\hat{I}_{a}$ from the real XYZ image $L_{a}$:
$\hat{I}_{a}=h(L_{a},\mathcal{E}_{a}).$ (2)
The $g(\cdot)$ and $h(\cdot)$ learn the inter-device general properties in ISP behaviors, while the device-aware embedding $\mathcal{E}_{a}$ focuses on device-specific intra-device properties of camera $a$.
The above formulation allows us to concurrently learn ISP behaviors of multiple cameras $\{a,b,c,...,z\}$ by training $g$ or $h$ alongside the device-aware embeddings $\{\mathcal{E}_{a},\mathcal{E}_{b},\mathcal{E}_{c},...,\mathcal{E}_{z}\}$.
3.2 Model Design
As depicted in Figure 2, our Uni-ISP utilizes the inverse ISP module $g$ and forward ISP module $h$, each featuring an encoder-decoder architecture. Both modules incorporate Local Feature Extraction Blocks (LFEBs) for detailed local processing and Global Feature Manipulation Blocks (GFMBs) for broad image adjustments, mirroring the dual processing dynamics of real camera ISPs that manage both global operations like exposure compensation and color correction, and local tasks such as tone mapping and highlight recovery.
Local Feature Extraction Blocks
Each encoder and decoder stage in Uni-ISP comprises four Local Feature Extraction Blocks (LFEBs). LFEBs in the encoder and decoder stages include max-pooling layers and upsampling layers respectively. Each LFEB contains multiple convolutional layers, activation layers, half instance normalization [6], and spatial/channel attention layers. Residual connections link LFEBs across the encoder and decoder stages.
Global Feature Manipulation Blocks
Global ISP operations are significantly affected by parameters such as exposure time and ISO, which alter the overall appearance of the photos. Our GFMBs are designed to incorporate these parameters to modify the residual features from the encoder’s LFEBs. These manipulated features are then relayed to corresponding LFEBs in the decoder stage. Used camera parameters (exposure, ISO, and f-number) are extracted from the EXIF metadata of JPEG images produced by cameras.
Device-aware Embedding Interaction Module
Positioned between the encoder and decoder stages, the Device-aware Embedding Interaction Modules (DEIM) enhance the model’s ability to adapt to different camera devices by interacting with device-aware embeddings. Given a device-aware embedding $\mathcal{E}_{a}$, the DEIM applies an attention-based transform to the bottleneck features $B$ from the encoder stage and output $F_{a}$. This setup enables the model to adaptively learn the ISP behaviors of multiple cameras concurrently.
Figure 3: The illustration of frequency bias in dataset wrapped using optical flow method. The interpolation during the wrapping will make images look blurry compared to the original one, eliminating its high-frequency component.
3.3 Training Scheme
We craft a special training scheme for Uni-ISP, which consists of two types of training objectives, i.e., self-camera ISP objective and cross-camera ISP objective.
3.3.1 Self-Camera Training Objective
In the self-camera training objective, our Uni-ISP learns the inverse and forward ISP behavior of multiple cameras simultaneously.
Given a sRGB image $I_{a}$ and corresponding XYZ image $L_{a}$ taken with camera $a$, the inverse ISP training objective is to minimize inverse ISP loss $\mathcal{L}_{\text{Inv}}$, aka the difference between $L_{a}$ and the predicted $\hat{L}_{a}$ given in Equation 1:
$\mathcal{L}_{\text{Inv}}=\|L_{a}-\hat{L}_{a}\|_{1}.$ (3)
The forward ISP training objective requires the model to minimize the following forward ISP loss $\mathcal{L}_{\text{For}}$ that calculates differences between the ground truth $I_{a}$ and the $\hat{I}_{a}$ given in Equation 2:
$\mathcal{L}_{\text{For}}=\|I_{a}-\hat{I}_{a}\|_{1}.$ (4)
3.3.2 Cross-Camera Training Objective
The target photo in the cross-camera training objective is taken by a different camera than the one that captures the input photo. Specifically, this training objective is applied on the forward ISP module $h$ to handle the camera model transition. Assuming $I_{a}$ is taken by camera model $a$ and $I_{b}$ is taken by camera model $b$, the inverse ISP module $g$ of Uni-ISP takes the $I_{a}$ as input and predicts the XYZ image $\hat{L}_{a}$, which is the same process as the one described in Equation 1. Then the forward ISP module $h$ convert the predicted $\hat{L}_{a}$ into the sRGB image $\hat{I}_{b}$ of camera $b$:
$\hat{I}_{b}=h(\hat{L}_{a},\mathcal{E}_{b}),$ (5)
where $\mathcal{E}_{b}$ indicate the device-aware embeddings of camera $b$. Here the $h$ takes the $\hat{L}_{a}$ input as input, which is different from the ground truth $L_{a}$ used in the self-camera training objective. The cross-camera training objective is defined to minimize the distance between $\hat{I}_{b}$ and the ground truth $I_{b}$.
In the above definition, the input image $I_{a}$ and the output ground truth $I_{b}$ can not be aligned since they are captured using different cameras. This makes pixel-level losses like L1-loss fail to drive the cross-camera training objective that minimizes the distance between $\hat{I}_{b}$ and $I_{b}$. Therefore, we first use the optical-flow-based method, RAFT [34], to wrap our dataset for the cross-camera training objective. All sRGB images $I_{b}$ taken with camera $b$ will be wrapped to be $I^{w}_{b}$, aligned with sRGB images $I_{a}$ taken with camera $a$. We apply this wrapping to every possible camera-to-camera sRGB pair in our dataset. Any areas that cannot be aligned will be annotated as occlusion and masked during training.
However, although the optical-flow-based wrapping method effectively aligns these images, it also introduces a frequency bias in our dataset. As demonstrated in Figure 3, the wrapped image tends to have less high-frequency details compared to the image before wrapping. If we use the aligned images to train our model directly, it will unexpectedly learn to smooth the images in the cross-camera ISP tasks.
To tackle this problem, we propose the frequency bias correction (FBC) loss for the cross-camera training objective. The Equation 9 shows the process of this task. The FBC loss can be written as:
$\mathcal{L}_{FBC}=\|f_{low}(\hat{I_{b}})-f_{low}(I^{w}_{b})\|_{1}+\mathcal{L}_{Freq}(\hat{I_{b}},I_{b}),$ (6)
where $f_{low}$ is a low-pass filter and $\mathcal{L}_{Freq}$ is a loss in frequency domain. In our implementation, the $f_{low}$ is a Gaussian filter with a kernel size of 5, and we adopt the focal frequency loss [15] as the $\mathcal{L}_{Freq}$.
Finally, we use the FBC loss $\mathcal{L}_{FBC}$ to drive the cross-camera training. Detailed ablation results on this loss function are presented in Section 4.5.
Figure 4: The preview of 3 scenes in our new dataset (left) and our capture devices (right). Each scene includes synchronized sRGB-Raw pairs of five smartphone cameras: Apple iPhone 14 Pro Max, Google Pixel 6 Pro, Huawei P40, Samsung Galaxy S20, and Xiaomi Mi 12. The raw images are visualized as XYZ images here, which can be converted back to raw without loss.
3.3.3 Overall Loss
In summary, during the training of Uni-ISP, we have three loss terms, the inverse ISP loss $\mathcal{L}_{\text{Inv}}$, the forward ISP loss $\mathcal{L}_{\text{For}}$, and the FBC loss $\mathcal{L}_{\text{FBC}}$ in total, and an additional regularization term $\mathcal{L}_{\text{NRR}}$:
$\mathcal{L}=\mathcal{L}_{\text{Inv}}+\mathcal{L}_{\text{For}}+\mathcal{L}_{\text{FBC}}+\lambda\mathcal{L}_{\text{NRR}},$ (7)
where $\mathcal{L}_{\text{NRR}}$ is the additional neutral rendering regularization and $\lambda$ is the balance weight for it. The neutral rendering regularization guides the model to learn a virtual camera that performs standard color conversion between XYZ and sRGB color space when the device-aware embedding is given as a zero vector $\emptyset$:
$\mathcal{L}_{\text{NRR}}=\|s(I_{a})-g(I_{a},\emptyset)\|_{1}+\|s^{-1}(L_{a})-h(L_{a},\emptyset)\|_{1},$ (8)
where $s(\cdot)$ and $s^{-1}(\cdot)$ is the sRGB-XYZ and XYZ-sRGB color conversion respectively. The neutral rendering regularization adds an anchor for users if they want to enhance or reduce the photographic appearance of a certain camera without inter/extrapolating with the device-ware embedding of another camera.
3.4 Novel Dataset
Although existing sRGB-RAW datasets allow for training models on both inverse and forward ISP tasks, there remains a need for datasets that contain sRGB-RAW pairs captured synchronously by multiple devices. Such datasets are essential for training models to handle cross-camera ISP tasks effectively, which are crucial for applications like photographic appearance transfer and inter/extrapolation.
To address this challenge, we collected a novel dataset named FiveCam, featuring synchronously captured sRGB-RAW pairs from five distinct camera models. This dataset encompasses 2,464 high-resolution (4K) raw and JPEG images representing approximately 500 diverse scenes. Cameras used in the FiveCam dataset include the Apple iPhone 14 Pro Max, Google Pixel 6 Pro, Huawei P40, Samsung Galaxy S20, and Xiaomi Mi 12. All cameras are synchronized using a programmed Bluetooth shutter to ensure consistent timing across all devices.
A preview of the FiveCam dataset, illustrated in Figure 4, showcasing its three scenes alongside the capture devices used. Scenes of our FiveCam dataset are richly varied, capturing both natural landscapes and urban environments under multiple lighting conditions, ranging from broad daylight to nighttime settings, and including both outdoor and indoor illuminance.
Additionally, we have created an sRGB-XYZ version of this dataset, where raw images from all cameras are processed using as-shot white balance, a linear demosaicing algorithm, and converted to the standard camera-to-XYZ color space. The XYZ images in this version retain their linearity, making them particularly beneficial for downstream tasks that require maintaining the raw image’s linearity.
This dataset serves dual purposes: as a conventional raw image dataset for learning the inverse and forward ISP, and as a specialized resource for training models on cross-camera ISP tasks like photographic appearance transfer and inter/extrapolation. Please refer to the supplementary material provided for more detailed information about this dataset.
4 Applications
Table 1: Quantitative results of inverse and forward ISP tasks in the multi-camera mixed test set. All models are trained and tested with data from five different cameras. Please refer to Section 3.4 for detailed model names for each camera. To ensure fairness, the CycleISP [40] is not included in the forward ISP test since it requires ground truth RGB input during the forward ISP process.
Method Inverse ISP Forward ISP
PSNR ($\uparrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$)
Cycle ISP [40] 28.836 0.8632 - -
CIE-XYZ Net [1] 24.990 0.7960 22.515 0.8750
InvISP [38] 26.380 0.8042 21.644 0.8625
ParamISP [18] 31.212 0.9180 26.739 0.9182
Uni-ISP (Ours) 32.699 0.9396 29.154 0.9307
In this section, we conduct four experiments to validate Uni-ISP on different applications, including 1) inverse and forward ISP, 2) photographic appearance transfer, 3) photographic appearance inter/extrapolation, and 4) zero-shot source camera forensics. We also conduct ablation studies and analytical experiments to verify the effectiveness of Uni-ISP.
Figure 5: Two scenes that compare our Uni-ISP with other methods in the task of inverse ISP. The difference maps between the prediction of each model and the ground truth are shown in the second and fourth rows. For better visualization, the XYZ images are adjusted with a 50% increase in brightness to make the content easier to observe.
Figure 6: Two scenes that compare our Uni-ISP with other methods in the task of forward ISP. The CycleISP [40] is not included for a fair comparison since its forward module requires information from the ground truth RGB image.
Figure 7: Quantitative results of inverse and forward ISP tasks in the single-camera test set. CycleISP [40], CIE-XYZ Net [1], InvISP [38], and ParamISP [18] are trained for each camera model separately while our device-aware Uni-ISP is trained on the mixed dataset. Numbers of total trained parameters for each method in this experiment are noted under the model name in gray. All models are tested with data from a single camera model (name mentioned in the horizontal axis), and the PSNR values are indicated in the vertical axis. For full numeric results in a table, please refer to the supplementary materials.
4.1 Inverse and Forward ISP
Settings.
We compare Uni-ISP with previous ISP learning methods, including CycleISP [40], CIE-XYZNet [1], InvISP [38], and ParamISP [18]. All methods are retrained on the FiveCam dataset and we train two versions of baseline methods. The first version is trained on the mixed dataset, where sRGB-XYZ image pairs of all five cameras are used for training. The second version is trained on every single camera separately. Our Uni-ISP only trains a single version on the mixed dataset, with solely the self-camera training objective.
Results.
For the first version, we show quantitative results in Table 1, and qualitative results in Figure 5 and Figure 6. By observing the images and difference maps, our Uni-ISP has less error compared to other methods in both two tasks, which is consistent with the numeric results shown in Table 1. For the second version, results in Figure 7 demonstrate our Uni-ISP still outperforms all previous methods when previous methods are trained on a single device separately, i.e., five models for each previous method. In contrast, for our Uni-ISP, we only need to train a single model.
Figure 8: Comparison among Uni-ISP, HDRnet [13], NILUT [8] in photographic appearance transfer. All three models are asked to transfer the photographic appearance to the camera that takes the ground truth photo. Our Uni-ISP shows more accurate transfer results.
4.2 Photographic Appearance Transfer
Settings.
Photographic appearance transfer from camera model $a$ to camera model $b$ can be defined as,
$I_{b}=h(g(I_{a},\mathcal{E}_{a}),\mathcal{E}_{b}),$ (9)
where $\mathcal{E}_{a}$ and $\mathcal{E}_{b}$ indicate two distinct device-aware embeddings of two camera models. This process uses inverse module $g$ to inverse the image to XYZ space and then uses forward module $h$ to generate the final image $I_{b}$. Given the absence of prior work, we select a series of global and local color transform methods as baselines, including learnable global color transforms, learnable 3D LUTs (Look-Up Tables), learnable bilateral grids, HDRnet [13] and NILUT [8]. We input the camera label as the style code within HDRnet and NILUT for a fair comparison. For the first three methods, we train 25 sub-models to accommodate the various transfer mappings required by the 25 possible pairs from the five devices in FiveCam. Still, we only need to train a single Uni-ISP. We adopt PSNR and DISTS [11] for evaluation.
Results.
As shown in Table 2, our method achieves the best results in terms of the two metrics, demonstrating the effectiveness of our method. A visual comparison presented in Figure 8 further illustrates that our Uni-ISP achieves results much closer to the ground truth compared to previous methods like HDRnet and NILUT, indicating the superior performance of our model. Moreover, our model has good temporal stability without a specific design. Figure 9 displays frames from the original iPhone video alongside those transferred to emulate the photographic styles of Google and Xiaomi. Please refer to the supplementary materials for the video.
Figure 9: Results of photographic appearance transfer on videos with Uni-ISP. We can see the consistent photographic appearance in the sky, even if our model does not have any temporal-specific design.
Table 2: Quantitative results of photographic appearance transfer. The best value of each column is highlighted in bold.
Method DISTS($\downarrow$) PSNR($\uparrow$)
Learned Color Matrices 0.1838 20.792
Learned 3D LUTs 0.2027 20.853
Learned Bilateral Grids 0.1971 20.625
HDRnet [13] 0.1722 20.863
NILUT [8] 0.1434 21.390
Uni-ISP (Ours) 0.1392 24.237
Table 3: Source camera identification accuracy of CMIResNet [7], MSFFN [14], and our Uni-ISP.
Method Accuracy
CMIResNet [7] 0.3833
MSFFN [14] 0.4333
Uni-ISP (Zero-shot) 0.8167
Figure 10: Photographic style interpolation and extrapolation between different camera models. We achieve this by adjusting the coefficient $\alpha$ in Eq. 10. Please refer to the electronic version on a bright display for better visualization.
4.3 Photographic Appearance Inter/Extrapolation
Settings.
In addition to transferring the photographic appearance, Uni-ISP presents the ability to inter/extrapolate the photographic appearance between two cameras smoothly. Given the photo $I_{a}$ captured by camera $a$, $I_{a}$ is firstly inverse to XYZ image by $g$. Then, the XYZ image is encoded to be the bottleneck feature and separately interacts with device-aware embeddings $\mathcal{E}_{a}$ and $\mathcal{E}_{b}$, producing intermediate features $F_{a}$ and $F_{b}$ right after the DEIM. Then, the manipulated photo $I_{a+b}$ can be obtained by:
$I_{a+b}=h_{decoder}((1-\alpha){F}_{a}+\alpha{F}_{b}),$ (10)
where $h_{decoder}$ is the decoder part of $h$, $\alpha\in[0,1]$ is the linear interpolation weight. For extrapolation, we relax the convex combination to the affine combination, where $\alpha$ is not limited in $[0,1]$.
Results.
The results of interpolating and extrapolating photographic appearance are demonstrated in Figure 10. In the first scenario, the Samsung Galaxy S20 prefers a deeper shadow and the extrapolation leads to a more “Samsung” style image. The second scenario is a night scene. Since Apple smartphones are famous for their night photography mode, our model also produces brighter results when we ask the model to be more “Apple”, where the lake and the trees become clearer.
4.4 Zero-shot Image Forensics
We perform two forensics tasks: source camera identification and image slice detection, without training on these tasks.
4.4.1 Source Camera Identification
Settings
During the inference of source camera identification, Uni-ISP performs the inverse and forward ISP tasks multiple times with the device-aware embeddings of candidate camera models. Then we measure the edit distance $D$ of each inference, and the predicted camera model is the one that leads to the minimal edit distance. We empirically choose the structural similarity (SSIM) [36] as the metric for evaluating editing distance $D$. We compare our model with the camera model identification residual neural network (CMIResNet) [7] and the multi-scale feature fusion network (MSFFN) [14]. These methods are trained on the same training set as the Uni-ISP model, whose training objectives are classification tasks on the five cameras. For testing, we collect an extra test set containing 60 sRGB photos taken with the same five cameras in our proposed dataset, but in different scenes.
Results.
Table 3 shows the numeric results of the source camera identification experiments on our dataset. Based on these results, our methods show higher accuracy over the classification-based methods for source camera identification, revealing a new possible perspective on this task.
4.4.2 Image Splice Detection
Settings
Our Uni-ISP offers additional capability in image splice detection. We can slightly modify its inference process for source camera identification to support image splice detection without extra training. Suppose an image $I_{a}$ is known to be taken with camera $a$, but it is altered. We can predict a map to indicate the areas likely to be altered using the self-consistency of inverse and forward ISP learned by Uni-ISP. This can be viewed as a pixel-level task in image forensics. We demonstrate this capability to show the versatility of our Uni-ISP.
Results.
Figure 11 demonstrates the example of Uni-ISP’s ability in image slice detection, where we put some fake objects on photos taken with iPhone 14 Pro Max. We harmonize the pasted fake objects using Poisson image editing [28]. The detection results are the visualization of the SSIM map, where white areas refer to minor differences, i.e., small editing distance, and darker areas indicate the location that may not belong to the original photo. The results prove the versatile ability when we unify the learning process of multiple camera ISPs.
Figure 11: Visualization of image slice detection by Uni-ISP. A fake object is integrated into a real photo using Poisson image editing [28], with darker regions in the result indicating suspected modifications.
4.5 Analysis
Evaluation on HDR Rendering
Learning the inverse and forward ISP benefits HDR rendering. To validate the HDR rendering quality with currently learned inverse and forward ISP models, we synthesize multi-exposure raw image stacks with digital gains of $[1,2.6,4.2]$ on our test set, render them into sRGB images with a fixed ISP pipeline, and use the exposure fusion to produce the HDR tone-mapped images as ground truths. Given an LDR sRGB image, we adopt models for inverse ISP to convert the LDR sRGB image into an XYZ image. Then, we synthesize a stack of frames with the same digital gains as the dataset and follow the same manner of the dataset to render them into an HDR tone-mapped image.
Figure 12 shows the visualized results and each model’s performance in PSNR on the HDR rendering task. Consistent with the inverse ISP performance shown in Table 1, our method also achieves the best quality in the HDR rendering task.
Figure 12: Visualization of the HDR rendering results and the average PSNR on the test set. Thanks to the accuracy in inverse ISP of Uni-ISP, it produces the closest HDR rendering results to the ground truth.
Ablation Study on FBC Loss
We conduct ablation on FBC loss by replacing it with an L1 loss that minimizes differences between model predictions and wrapped images for training. As shown in the image and spectrum visualization of Figure 13, the model trained with FBC loss produces clearer results than the one trained without FBC loss. Numeric results are presented in the supplementary material, which also demonstrates the FBC loss effectively avoids the frequency bias in the wrapped dataset.
Figure 13: Ablation results on the FBC loss. With FBC loss, we effectively avoid our model to learn the frequency bias in our wrapped dataset. The spectrum visualization of the noted patch also demonstrates that the model trained with FBC loss produces a sharper image with stronger high-frequent components.
Visualization of Internal Features
To elucidate why Uni-ISP effectively handles seamless inter/extrapolation of photographic appearances across multiple cameras, we visualize the internal feature distributions of our model. Specifically, we execute both self-camera tasks (inverse and forward ISP) and cross-camera tasks (photographic appearance transfer) using our test set. During these tests, we capture the forward ISP module $h$’s intermediate bottleneck features $B$ before they are processed by the DEIM and $F_{x}$ for each camera $x\in\{a,b,c,d,e\}$ after they are processed by the DEIM. Here, $x\in\{a,b,c,d,e\}$ denotes the variable for the five cameras in our new dataset. The location of $B$ and $F_{x}$ are noted in Figure 2. We utilize UMAP [26] to project these high-dimensional features into a 2D space, as shown in Figure 14. This visualization clearly delineates the transformation dynamics of features, with distinct clusters emerging for each camera model. This clustering visually confirms that the process of interpolating or extrapolating photographic appearances mimics sampling from a well-defined manifold of camera ISP behaviors. Thanks to the DEIM, which transforms features from an indistinct subspace (points noted by triangles) into a clearly delineable manifold (points noted by circles and squares), our model facilitates smooth inter/extrapolation between the photographic appearances of different camera models, as depicted in Figure 10.
Figure 14: UMAP visualization of the internal features $B$ right before the DEIM and internal features $F_{x}$ ($x\in\{a,b,c,d,e\}$) right after the DEIM inside the forward ISP module $h$ of Uni-ISP. The DEIM lets corresponding device-aware embeddings $\mathcal{E}_{x}$ interact with the bottleneck features produced by the encoder stage of forward ISP module ($h_{encoder}$), then produce the $F_{x}$. Features $F_{x}$ lie in a special manifold, where features in both self-camera and cross-camera tasks with the same target camera models are clustered together.
Few-shot Extension Ability
Given a few sRGB-Raw samples of an unseen camera, our Uni-ISP has the extension ability to perform well in inverse and forward ISP learning of that new camera by only training a new device-aware embedding while keeping the rest of the model frozen. We conduct the few-shot extension experiments on the S7 ISP dataset [32] to validate the few-shot extension ability of Uni-ISP. During training, we load the weights of the previously trained model in Section 4.1 and only optimize the newly created device-aware embedding for this unseen camera, Samsung Galaxy S7. For comparison, we select the ParamISP [18] since it outperforms previous methods in most cases. To simulate the scenario where only limited training data is available, we randomly select 10 scenes of the whole S7 dataset for training and the other 100 scenes for testing the generalizability. All models are trained for 50 epochs for convergence. Tests are conducted on the checkpoint of the 50th epoch. Table 4 shows the results of few-shot extension experiments with only 10 samples for learning this unseen camera model’s inverse and forward ISP.
As we can see, our Uni-ISP outperforms the ParamISP [18] in forward ISP when only 10 training samples are provided, which demonstrates the value of this extension ability in extreme scenarios of lacking adequate training datasets.
Table 4: Few-shot extension experiments on S7 ISP dataset [32] with only 10 samples for learning this unseen camera model’s inverse and forward ISP. We compare performances of the extension version of Uni-ISP (Uni-ISP-Ext) with the trained-from-scratch version of ParamISP [18] since the ParamISP outperforms other prior models in most cases. Numbers of trained parameters are given in the Param. column. The best performance for each few-shot sample scenario is noted in bold.
Method Param. Inverse ISP Forward ISP
PSNR ($\uparrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$)
ParamISP 700K 23.907 0.8305 21.439 0.8489
Uni-ISP-Ext 256 25.307 0.8597 22.798 0.8590
5 Conclusion
In this study, we presented the Uni-ISP, which can concurrently learn inverse and forward ISP behaviors of multiple cameras. The Uni-ISP harnesses the commonalities and distinctions across different cameras, enhancing the performance of ISP tasks beyond that of models trained on single-camera or mixed datasets. Additionally, Uni-ISP introduces novel applications for learned ISPs, including photographic appearance editing and zero-shot camera source forensics.
In future work, we plan to explore the interaction between the device-specific components and the backbone network by introducing more domain knowledge from ISP designs. Furthermore, considering the complexity of inverse ISP tasks, which can often be highly ill-posed, investigating the use of generative priors to maintain image fidelity presents an intriguing avenue for future research.